World Population was 6.7 billion and the exponential growth rate was 1.14% per year. source : u.s. Census Bureau
A. Find the exponential growth function
B. Predict the world population in 2014
C. When will the world population be 8.0 billion
can you explain the steps using the formula (P(t)

Answer by stanbon(75887)

In 2008,the world population was 6.7 billion and the exponential growth rate was 1.14% per year.
A. Find the exponential growth function
A(t) = 6.7 bil (1.14)^t where t is the number of years after 2008.
----------------------------------
B. Predict the world population in 2014
A(6) = 6.7 bil (1.14)^6 = 6.7 bil (2.1950) = 14.71 billion
-------------------------
C. When will the world population be 8.0 billion
8 bil = 6.7 bil (1.14)^t
---
1.194 = 1.14^t
---
Take the log:
t*log(1.14) = log(1.194)
t = 1.35 years
=====================
Cheers,
Stan H.